OpenStudy (anonymous):

Help please! Question in comments:

OpenStudy (anonymous):

Terri is rewriting \[7^{13}\] in expanded form. How many factors should she use in her expression?

OpenStudy (australopithecus):

\[x^3 = x*x*x\] has 3 factors \[x^4 = x*x*x*x\] has 4 factors \[x^n=x*x*x....\] has n factors